# Can Knowledge Graphs Make Large Language Models More Trustworthy? An Empirical Study Over Open-ended Question Answering

### v3

@@ A.1 Query Construction
In this section, we discuss the details of the query construction of OKGQA. We first introduce the human-in-the-loop process to optimize the instruction for generating the queries, as shown in §A.1.1. We then present the metrics for quantify the generated queries in §A.1.2. Subsequently, we provide experiments results of human-in-the-loop process and demonstrate the Pearson correlation coefficients between human and LLM scores across rounds of optimization, and verify the inter-rater reliability across our evaluators in §A.1.3.
@@ A.1.1 Human-in-the-loop for instruction optimization
To ensure that the generated queries accurately represent real-world scenarios and complexities, we propose a human-in-the-loop process to optimize the instruction used for generation, as shown in Figure 7. To ensure clarity, we summarize this optimization process here:
@@
The step 3 and 4 are conducted by prompting LLM with prompt specified in §D.3, and steps 1 to 4 are running iteratively to reducing $s_{\mathbf{auto}}$ and $s_{\mathbf{human}}$ discrepancies. This process quite mimics the way of reinforcement learning with human feedback (RLHF) (Ouyang et al., 2022) and inherits the benefit that labeling the reward of the LLMs’ output is much easier than labeling the output directly.
@@ A.1.3 Verifying human-in-the-loop
In addition, we also consider verifying the inter-rater reliability across three evaluators as shown in Table 4. We report the Cohen’s Kappa coefficient for each pair of evaluators as follows. Using the Landis and Koch (1977) interpretation guidelines, the Cohen’s Kappa coefficients for Naturalness and Relevance (ranging from 0.79 to 0.85) fall within the “Substantial” to “Almost Perfect” categories, indicating strong inter-rater reliability for these metrics. This reflects a shared understanding of the evaluation criteria, resulting in consistent ratings among evaluators. For Specificity, Novelty, and Actionability, the coefficients range from 0.58 to 0.68, placing them primarily in the “Moderate” to “Substantial” categories. These results suggest moderate reliability for these metrics, likely due to subjective interpretation and less clearly defined evaluation guidelines. Novelty, with lower coefficients around 0.61 to 0.63, highlights variability in ratings, suggesting that evaluators may have differing perspectives on what qualifies as novel (but the inter-rater reliability is still considered “Substantial”). Meanwhile, Actionability performs slightly better, nearing the “Substantial” range, indicating moderately consistent criteria.
@@ A.3 Prize-Cost-based Path Retrieval
In this section, we detail the path-retrieval method used in §3.2. It is designed to construct and evaluate paths in a graph based on predefined prize assignments and cost allocations. The objective is to form sequences of nodes and edges, represented as $\mathcal{P}=\{v_{1},e_{1},v_{2},\dots,e_{n-1},v_{n}\}$, that maximize the overall score and minimize the costs. To efficiently manage the exploration of potential paths, we utilize a priority queue, a data structure that allows paths to be organized based on their scores, ensuring that the highest-scoring paths are processed first. The method starts by picking a number of starting nodes with high prizes. We then expand these starting points by exploring neighboring nodes. For each neighbor, the method calculates a new score. This score is the sum of the neighbor’s prize and the edge’s prize minus the edge’s cost. If this neighbor hasn’t been visited before, which helps avoid looping, the algorithm adds this neighbor to the path. This new path is then added to the priority queue. This expansion keeps going until the path reaches a maximum length or can’t be extended further. The algorithm keeps track of paths already explored to avoid repetition and ensure paths don’t loop back on themselves. When no more paths can be added or the priority queue is empty, the algorithm sorts the paths by their scores from highest to lowest.
@@ A.4 LLM Evaluation Clarity
To address the concern regarding potential self-enhancement bias in LLM-as-evaluator frameworks, we have conducted extensive validation of our evaluation approach. In specific, we randomly sample 100 questions and evaluated them using three different LLMs (gpt-4o-mini, llama-3.1-8b-instruct, and gemma-2-9b-it). We measured inter-model agreement using Cohen’s Kappa as shown in Table 6, which showed substantial to almost perfect consistency. This indicates that the evaluation results are consistent across different LLMs, even when the model generating the responses is not the same as the one evaluating them (e.g., using gpt-4o-mini for generation and llama-3.1-8b-instruct for evaluation). These findings confirm that the evaluation is robust and independent of the specific LLM used as the evaluator.
@@
In addition, we also collect human evaluations for these 100 samples. Three experts annotators rate each anonymized response on context relevance, comprehensiveness, correctness and empowerment using a 1-5 Likert scale. The average human ratings are computed and compared with automated scores using G-Eval. The Pearson’s correlation yields a score of 0.78, indicating strong alignment between human judgment and LLM-based evaluation. Combined with the inter-model agreement shown in Table 6, these results confirm that our evaluation is robust, consistent, and largely independent of the specific LLM used as the evaluator.
@@ Appendix C Related Work
Due to the stochastic decoding process of Large Language Models (LLMs), i.e., sampling the next token in the sequence, LLMs exhibit probabilistic behaviors: (1) potentially yielding varied outputs of the same input across different instances (Agrawal et al., 2023); (2) cannot accurately interpret phrases or terms when the context is vague and resides in a knowledge gap of the model. This will lead to outputs that may sound plausible but are often irrelevant or incorrect. This will lead to outputs that may sound plausible but are often irrelevant or incorrect. This “hallucinations” undermines the reliability of LLMs (Huang et al., 2023). One emerging research trend is enhancing LLMs through integrating external knowledge graphs (Agrawal et al., 2023). KGs offer structured, explicit, and up-to-date factual knowledge, including domain-specific knowledge, providing a faithful knowledge source for reasoning (Zheng et al., 2023; Agrawal et al., 2023; Sui et al., 2022). Moreover, each piece of information in KGs can be traced back to its source, providing context and provenance. This traceability not only aids in verifying the reliability of the information but also provides clear pathways of reasoning, making the interpretation process transparent.
Researchers employ diverse strategies to augment the LLMs by integrating external KGs (Sui et al., 2024; He et al., 2024b). For example, KAPING (Baek et al., 2023b) matches entities in questions to retrieve related triples from knowledge graphs for zero-shot question answering. Wu et al. (2023) finds that converting these triples into textualized statements can further enhance LLM performance. StructGPT (Jiang et al., 2023b) propose to convert user query into structured formats (e.g., SPARQL) for information extraction from KGs. Following the succuess of internal reasoning-enhancement methods like Chain-of-thoughts (CoT) (Wei et al., 2022), Reflexion (Shinn et al., 2024), and Tree-of-thoughts (ToT), He et al. (2022) propose “rethinking with retrieval” to use decomposed reasoning steps from CoT prompting to retrieve external knowledge, leading to more accurate and faithful explanations. IR-CoT (Trivedi et al., 2022b) interleaves the generation of CoT with knowledge retrieval from corresponding KGs, iteratively guiding both retrieval and reasoning for multi-step questions. MindMap (Wen et al., 2023) introduce a plug-and-play approach to evoke graph-of-thoughts reasoning in LLMs. Similarly, RoG (Luo et al., 2023b) use KGs to create faithful reasoning paths based on various relations, enabling interpretable reasoning in LLMs.

### v4

@@ A.1 Query Construction
In this section, we discuss the details of the query construction of OKGQA. We first introduce the human-in-the-loop process to optimize the seed instruction for generating the queries, as shown in §A.1.1. We then present the metrics to quantify the generated queries in §A.1.2. Subsequently, we provide experiments results of human-in-the-loop process and demonstrate the Pearson correlation coefficients between human and LLM scores across different rounds of optimization, and verify the inter-rater reliability across different LLM evaluators in §A.1.3.
@@ A.1.1 Human-in-the-loop for instruction optimization
To ensure that the generated queries represent real-world scenarios and complexities, we propose a human-in-the-loop process to optimize the seed instruction used for generation, as shown in Figure 7. To ensure clarity, we summarize this optimization process here:
@@
Steps 3 and 4 are performed by prompting the LLM with the instruction specified in §D.3, and the entire process from steps 1 to 4 is iterated to minimize the discrepancy between $s_{\mathbf{auto}}$ and $s_{\mathbf{human}}$. This procedure closely resembles the way of reinforcement learning with human feedback (RLHF) (Ouyang et al., 2022) and inherits the benefit that labeling the reward of the LLMs’ output is much easier than directly labeling the outputs.
@@ A.1.3 Verifying human-in-the-loop
In addition, we also consider verifying the inter-rater reliability across three evaluators as shown in Table 4. We report the Cohen’s Kappa coefficient for each pair of evaluators. According to the interpretation guidelines from Landis and Koch (1977), the Cohen’s Kappa coefficients for Naturalness and Relevance (ranging from 0.79 to 0.85) fall within the “Substantial” to “Almost Perfect” categories, indicating strong inter-rater reliability for these metrics. This indicates that the evaluators share a common understanding of the evaluation criteria, leading to consistent ratings across evaluators. For Specificity, Novelty, and Actionability, the coefficients range from 0.58 to 0.68, placing them primarily in the “Moderate” to “Substantial” categories. These results suggest moderate reliability for these metrics, likely due to subjective interpretation and less clearly defined evaluation guidelines. Novelty, with lower coefficients around 0.61 to 0.63, highlights variability in ratings, suggesting that evaluators may have differing perspectives on what qualifies as novel (but the inter-rater reliability is still considered “Substantial”). Meanwhile, Actionability performs slightly better, nearing the “Substantial” range, indicating moderately consistent criteria.
@@ A.3 Prize-Cost-based Path Retrieval
In this section, we detail the path-retrieval method used in §3.2. It is designed to construct and evaluate paths in a graph based on predefined prize assignments and cost allocations. The objective is to form sequences of nodes and edges, represented as $\mathcal{P}=\{v_{1},e_{1},v_{2},\dots,e_{n-1},v_{n}\}$, that maximize the overall score and minimize the costs. To efficiently manage the exploration of potential paths, we utilize a priority queue, a data structure that allows paths to be organized based on their scores, ensuring that the highest-scoring paths are processed first. The method starts by picking a number of starting nodes with high prizes.From each starting node, it expands by exploring neighboring nodes. For each neighbor, a new score is computed as the sum of the neighbor’s prize and the edge’s prize minus the edge’s cost. If this neighbor has not been visited before, the algorithm appends it to the current path and adds this extended path to the priority queue. This expansion process continues until paths reach a maximum length or can no longer be extended. The algorithm maintains a record of explored paths to avoid repetition and cycles. Once no further expansions are possible or the priority queue is empty, the algorithm sorts the collected paths in descending order of their scores.
@@ A.4 LLM Evaluation Clarity
To address the concern regarding potential self-enhancement bias in LLM-as-evaluator frameworks, we provide extensive validation of our evaluation approach. In specific, we randomly sample 100 questions and evaluated them using three different LLMs (i.e., gpt-4o-mini, llama-3.1-8b-instruct, and gemma-2-9b-it). We measure the inter-model agreement using Cohen’s Kappa as shown in Table 6. It indicates that the evaluation results are consistent across different LLMs, even when the model generating the responses is not the same as the one evaluating them (e.g., using gpt-4o-mini for generation and llama-3.1-8b-instruct for evaluation). These findings confirm that the evaluation is robust and independent of the specific LLM used as the evaluator. In addition, we also collect human evaluations for these 100 samples. Three expert annotators rate each anonymized response on context relevance, comprehensiveness, correctness, and empowerment using a 1–5 Likert scale. The average human ratings are then computed and compared with automated scores obtained via G-Eval. The Pearson correlation coefficient of 0.78 indicates strong alignment between human judgments and LLM-based evaluations. Together with the inter-model agreement reported in Table 6, these results demonstrate that our evaluation is robust, consistent, and largely independent of the specific LLM used as the evaluator.
@@ Appendix C Related Work
Due to the stochastic decoding process of Large Language Models (LLMs), i.e., sampling the next token in the sequence, LLMs exhibit probabilistic behaviors: (1) potentially yielding varied outputs of the same input across different instances (Agrawal et al., 2023); (2) cannot accurately interpret phrases or terms when the context is vague and resides in a knowledge gap of the model. This will lead to outputs that may sound plausible but are often irrelevant or incorrect. This will lead to outputs that may sound plausible but are often irrelevant or incorrect. This “hallucinations” undermines the reliability of LLMs (Huang et al., 2023). One emerging research trend is enhancing LLMs through integrating external knowledge graphs (Agrawal et al., 2023). KGs offer structured, explicit, and up-to-date factual knowledge, including domain-specific knowledge, providing a faithful knowledge source for reasoning (Sui, 2021; Zheng et al., 2023; Agrawal et al., 2023). Moreover, each piece of information in KGs can be traced back to its source, providing context and provenance. This traceability not only aids in verifying the reliability of the information but also provides clear pathways of reasoning.
Researchers employ diverse strategies to augment the LLMs by integrating external KGs (Sui et al., 2024; He et al., 20 2, 2025; Sui et al., 2022). For example, KAPING (Baek et al., 2023b) matches entities in questions to retrieve related triples from knowledge graphs for zero-shot question answering. Wu et al. (2023) finds that converting these triples into textualized statements can further enhance LLM performance. StructGPT (Jiang et al., 2023b) propose to convert user query into structured formats (e.g., SPARQL) for information extraction from KGs. Following the succuess of internal reasoning-enhancement methods like Chain-of-thoughts (CoT) (Wei et al., 2022), Reflexion (Shinn et al., 2024), and Tree-of-thoughts (ToT), He et al. (2023a) propose “rethinking with retrieval” to use decomposed reasoning steps from CoT prompting to retrieve external knowledge, leading to more accurate and faithful explanations. IR-CoT (Trivedi et al., 2022b) interleaves the generation of CoT with knowledge retrieval from corresponding KGs, iteratively guiding both retrieval and reasoning for multi-step questions. MindMap (Wen et al., 2023) introduce a plug-and-play approach to evoke graph-of-thoughts reasoning in LLMs. Similarly, RoG (Luo et al., 2023b) use KGs to create faithful reasoning paths based on various relations, enabling interpretable reasoning.